Determine the [OH−] , pH, and pOH of a solution with a [H+] of 3.0×10−8 M at 25 °C.

[OH−]= 3.33x10-7 M

pH= 7.52

pOH=6.48

Determine the [H+], pH, and pOH of a solution with an [OH−] of 5.8×10−9 M at 25 °C.

[H+]= 1.72x10-6 m

pH= ?

pOH= ?

Determine the [H+][H+] , [OH−][OH−] , and pOH of a solution with a pH of 7.92 at 25 °C.

[H+]= ?

[OH−]= ?

pOH= 6.08

Determine the [H+] , [OH−] , and pH of a solution with a pOH of 11.00 at 25 °C.

[H+]= ?

[OH−]= ?

pH= 3
Answer #1

Part a

[H+] = 3.0*10^-8 M

pH = - log [H+] = - log (3.0*10^-8)

= 7.52

pOH = 14 - pH = 14 - 7.52 = 6.48

[OH-] = 10^-pOH = 10^-(6.48) = 3.31*10^-7 M

Part b

[OH-] = 5.8*10-9 M

pOH = - log [OH-] = - log (5.8*10^-9) = 8.24

pH = 14 - pOH = 14 - 8.24 = 5.76

[H+] = 10^-pH = 10^-5.76 = 1.74*10^-6 M

Part c

pH = 7.92

pOH = 14 - pH = 14 - 7.92 = 6.08

[H+] = 10^-pH = 10^-7.92 = 1.20*10^-8 M

[OH-] = 10^-pOH = 10^-6.08 = 8.32*10^-7 M

Part d

pOH = 11.00

pH = 14 - pOH = 14 - 11 = 3.00

[H+] = 10^-pH = 10^-11 M

[OH-] = 10^-pOH = 10^-3 M